OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 2 3 All real numbers No solution

hartnn (hartnn):

Multiply the brackets out: a(b+c) = ab+ac so whats -2(x+3) = ?

OpenStudy (anonymous):

all real numbers

OpenStudy (anonymous):

I got it, thanks :D